# On the Isbell Problem

### v1

@@ II $\tau_{0}(f\upharpoonright 1)$ $\tau_{0}(f\upharpoonright 2)$ $\tau_{0}(f\upharpoonright 3)$ldots
  5. 4)*
    $\Gamma_{\beta}^{0}=\Gamma_{<\beta+1}^{0}[\Gamma_{<\beta}^{0}]=\{p(\beta)[\Gamma_{<\beta}^{0}]:p\in\Gamma_{\beta+1}^{0}\}$.
  6. 5)
    $\Gamma_{0}^{1}=\Gamma_{<1}[\Gamma_{0}^{0}]=\{q[\Gamma_{0}^{0}]:(p,q)\in\Gamma_{<1}\}$.
  7. 6)
    $\Gamma_{\beta}=\Gamma_{<\beta+1}[\Gamma_{<\beta}^{0}]$.
  8. 7)
    $\Gamma_{\beta}^{1}=\Gamma_{\beta}[\Gamma_{<\beta+1}^{0}]$
  9. 8)
    $\Gamma_{<\beta}^{1}=\{\langle q(\beta)[\Gamma_{\beta+1}^{0}]:\beta\in dom(q)\rangle:(\exists p)(p,q)\in\Gamma_{<\beta}\}$
@@
Let $\mathbf{P}=\langle\mathbb{P}_{\beta},(\dot{\mathbb{Q}}_{\beta}^{0},\dot{\mathbb{Q}}_{\beta}^{1}):\beta<\alpha\rangle$ be a $\varsigma$-iteration and $\beta_{0}\leq\alpha$ some fixed ordinal. Define $\mathbf{P}^{(*\beta)}$ as:
@@
and the order given by $(p_{1},q_{1})\leq(p_{0},q_{0})$ if and only if $p_{1}\leq_{\mathbf{P}_{0}}p_{0}$ and for each $\gamma\in dom(q_{0})$, $p_{1}\upharpoonright\gamma\Vdash q_{1}(\gamma)\leq q_{0}(\gamma)$.
Note that for $\beta_{0}=\alpha$, for any $(p,q)\in\mathbf{P}^{(*\beta_{0})}$, we have $q=\emptyset$, so we identify $\mathbf{P}^{(*\alpha)}$ with $\mathbf{P}^{\alpha}_{0}=\mathbb{P}_{\alpha}$.
@@
Let $\mathbf{P}$ be a $\varsigma$-iteration. Then, for any $\beta<\alpha$, $\mathbf{P}$ is equivalent to $\mathbf{P}^{\beta}*[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$.
@@
Define $\Psi_{\beta}:\mathbf{P}\to\mathbf{P}^{\beta}*[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$ by
$\Psi_{\beta}(p,q)=((p\upharpoonright\beta,q\upharpoonright\beta),(p,q*\beta))$
We claim that $\Psi_{\beta}$ is a dense embedding. Fix $(r,\dot{s})\in\mathbf{P}^{\beta}*[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$. Extend $r$ to a condition $r_{1}$ which determines $\dot{s}$, that is, there is $s_{1}\in\mathbf{P}^{(*\beta)}$ such that $r_{1}\Vdash\dot{s}=s_{1}$. Assume $r_{1}=(p_{1},q_{1})$ and $s_{1}=(p_{1}^{\prime},q_{1}^{\prime})$. Therefore,
@@
    $(p_{1},q_{1})\Vdash p_{1}^{\prime}\upharpoonright\beta\in\Gamma_{<\beta}^{0}$.
@@
    $(r_{1},s_{1})\leq(r,\dot{s})$.
Since $(p_{1},q_{1})\Vdash p_{1}\upharpoonright\beta\in\Gamma_{<\beta}^{0}$, we have $p_{1}^{\prime}\upharpoonright\beta$ and $p_{1}$ are compatible. Let $p^{\prime\prime}_{2}$ be a common extension of $p_{1}$ and $p_{1}^{\prime}\upharpoonright\beta$, and define $p_{2}=p_{2}^{\prime\prime\frown}(p_{1}^{\prime}*\beta)$ and $q_{2}=q_{1}^{\frown}q_{1}^{\prime}$. Then, $(p_{2},q_{2})\in\mathbf{P}$, and
$\Psi_{\beta}(p_{2},q_{2})=((p_{2}^{\prime\prime},q_{1}),(p_{2}^{\prime\prime\frown}(p_{1}^{\prime}*\beta),q_{1}^{\prime}))\leq((p_{1},q_{1}),(p_{1}^{\prime},q_{1}^{\prime}))\leq(r,\dot{s})$
It is also easy to see that $(p_{1},q_{1})\leq(p_{0},q_{0})$ if and only if $\Psi_{\beta}(p_{1},q_{1})\leq\Psi_{\beta}(p_{0},q_{0})$. Thus, we have that $\Psi_{\beta}[\mathbf{P}]$ is a dense subset of $\mathbf{P}^{\beta}*[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$ isomorphic to $\mathbf{P}$. ∎
Notation. From now on, $\Psi_{\beta}:\mathbf{P}\to\mathbf{P}^{\beta}*[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$ will denote the function defined in the previous proof,
$\Psi_{\beta}(p,q)=(p\upharpoonright\beta,q\upharpoonright\beta,(p,q*\beta))$
Remark. As noted at the end of the proof of the previous lemma, $\Psi_{\beta}$ gives us a natural dense subset of $\mathbf{P}^{\beta}*[\mathbf{P}^{(*\beta)}:\mathbf{P}_{0}^{\beta}]$, to say
$\Psi_{\beta}[\mathbf{P}]=\{((p\upharpoonright\beta,q\upharpoonright\beta),(p,q*\beta)):(p,q)\in\mathbf{P}\}$
Note that $\Psi_{\beta}[\mathbf{P}]$ looks quite similar to a two step iteration, and it is equivalent to a two step iteration, but it is not a two step iteration. In what follows, we will identify $\mathbf{P}^{\beta}*[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$ with $\Psi_{\beta}[\mathbf{P}]$. This will free us from making some cumbersome definitions to ensure certain sets are open dense; however, it has the drawback of stating a few more necessary lemmas to justify the application the notion of $\varsigma$-properness of the pair $(\mathbf{P}^{\beta}_{0},\mathbf{P}^{\beta})$ to the almost $\mathbf{P}_{0}^{\beta}$-restricted iteration $\mathbf{P}^{\beta}\star[\mathbf{P}^{(*\beta)}:\mathbf{P}_{0}^{\beta}]$ (see below for a clarification of this assertion), since $\Psi_{\beta}[\mathbf{P}]$ is not exactly a two step iteration, even when it is forcing equivalent to a two step iteration. To differentiate $\Psi_{\beta}[\mathbf{P}]$ from $\mathbf{P}^{\beta}*[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$, while keeping the emphasis in the factorization, we denote $\Psi_{\beta}[\mathbf{P}]$ by $\mathbf{P}^{\beta}\star[\mathbf{P}^{(*\beta)}:\mathbf{P}_{0}^{\beta}]$.
Notation. Let $\mathbf{P}$ be a $\varsigma$-iteration of length $\alpha$ and $\beta<\alpha$.
@@
    $\mathbf{P}^{\beta}\star[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]=\Psi_{\beta}[\mathbf{P}]$.
@@
    To short the notation, if $\dot{\mathbb{R}}$ is a $\mathbf{P}^{\alpha}_{0}$-name, $\mathbf{P}^{\beta}\star[\mathbf{P}^{(*\beta)}*{\dot{\mathbb{R}}}:\mathbf{P}^{\beta}_{0}]$ denotes the forcing
    $\mathbf{P}^{\beta}\star[\mathbf{P}^{(*\beta)}*_{\mathbf{P}_{0}^{\alpha}}{\dot{\mathbb{R}}}:\mathbf{P}^{\beta}_{0}]$
    while $[\mathbf{P}^{(*\beta)}*\dot{\mathbb{R}}:\mathbf{P}^{\beta}_{0}]$ denotes the quotient forcing
    $[\mathbf{P}^{(*\beta)}*_{\dot{\mathbf{P}}_{0}^{\alpha}}\dot{\mathbb{R}}:\mathbf{P}^{\beta}_{0}]$
@@
    Similarly, $[\mathbf{P}*\dot{\mathbb{R}}:\mathbf{P}^{\beta}]$ denotes the quotient forcing
    $[\mathbf{P}*_{\dot{\mathbf{P}}_{0}^{\alpha}}\dot{\mathbb{R}}:\mathbf{P}^{\beta}]$
The context should be enough to avoid ambiguities.
Also note that for any generic filter $\Gamma_{<\beta}$, in $V[\Gamma_{<\beta}]$, we have that $[\mathbf{P}^{(*\beta)}:\mathbf{P}_{0}^{\beta}]=(\Psi_{\beta}[\mathbf{P}]^{\divideontimes})_{\Gamma_{<\beta}}$.
Now we explain what we mean that we have to justify the application of the notion $\varsigma$-properness of the pair $(\mathbf{P}^{\beta}_{0},\mathbf{P}^{\beta})$ to $\mathbf{P}^{\beta}\star[\mathbf{P}^{(*\beta)}:\mathbf{P}_{0}^{\beta}]$. A $\mathbf{P}^{\beta}_{0}$-restricted iteration of $\mathbf{P}^{\beta}$ and $[\mathbf{P}^{(*\beta)}:\mathbf{P}_{0}^{\beta}]$ is of the form $\mathbf{P}^{\beta}*_{\mathbf{P}^{\beta}_{0}}[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$ and conditions are ordered pairs $((p,q),\dot{r})$ where $(p,q)\in\mathbf{P}^{\beta}$ and $\dot{r}$ is a $\mathbf{P}^{\beta}_{0}$-name such that $p\Vdash_{\mathbf{P}_{0}^{\beta}}\dot{r}\in[\mathbf{P}^{(*\beta)}:\mathbf{P}_{0}^{\beta}]$. However, for any condition $(a,b)\in\mathbf{P}^{\beta}\star[\mathbf{P}^{(*\beta)}:\mathbf{P}_{0}^{\beta}]=\Psi_{\beta}[\mathbf{P}]$, we have that $a\in\mathbf{P}^{\beta}$, but $b$ is a ground model set, and not a $\mathbf{P}^{\beta}_{0}$-name. Thus, strictly speaking, when we have some family $\{D_{n}:n\in\omega\}$ of open dense subsets of $\mathbf{P}^{\beta}\star[\mathbf{P}^{(*\beta)}:\mathbf{P}_{0}^{\beta}]\cap\mathcal{M}$, where $\mathcal{M}$ is a suitable countable elementary submodel, the phrasing “$(p_{0},q_{0})\leq(p,q)$ and $(p_{0},q_{0})$ is $(\mathbf{P}_{0}^{\beta},\mathbf{P}^{\beta},\mathcal{M})$-generic for $\{D_{n}:n\in\omega\}$ and $[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$” does not completely fit into the $\varsigma$-proper definition. The intention of the Lemma 6.16 is to justify this use.
Lemma 6.16.
Let $\mathbf{P}$ be a $\varsigma$-iteration of length $\alpha$ and $\beta<\alpha$. Assume that $(\mathbf{P}_{0}^{\beta},\mathbf{P}^{\beta})$ is $\varsigma$-proper. Let $\mathcal{M}\prec H(\theta)$ be countable and such that $\mathbf{P},\beta\in\mathcal{M}$. Let $\{D_{n}:n\in\omega\}$ be a family of open dense subsets of $\mathbf{P}^{\beta}\star[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]\cap\mathcal{M}$. Then, for any $(p,q)\in\mathbf{P}^{\beta}\cap\mathcal{M}$, there is $(p_{0},q_{0})\in\mathbf{P}^{\beta}$ such that $(p_{0},q_{0})\leq(p,q)$ and a family of $\mathbf{P}_{0}^{\beta}$-names $\{\dot{E}_{n}:n\in\omega\}$ such that for each $n\in\omega$:
  1. $\circledast_{0}$
    $(p_{0},q_{0})$ forces $D_{n}^{\divideontimes}$ is an open dense subset of $[\mathbf{P}^{(*\beta)}:\mathbf{P}_{0}^{\beta}]\cap\mathcal{M}[\dot{\Gamma}_{<\beta}]$;
  2. $\circledast_{1}$
    $p_{0}$ forces each $\dot{E}_{n}$ to be an open dense subset of $[\mathbf{P}^{(*\beta)}:\mathbf{P}_{0}^{\beta}]\cap\mathcal{M}[\dot{\Gamma}_{<\beta}]$.
  3. $\circledast_{2}$
    $(p_{0},q_{0})\Vdash\dot{E}_{n}\subseteq D_{n}^{\divideontimes}$.
Thus, we extend the definition of $\varsigma$-properness to include posets of the form $\mathbf{P}^{\beta}\star[\mathbf{P}^{(*\beta)}:\mathbf{P}_{0}^{\beta}]$, we extend the definition of $(\mathbf{P}^{\beta}_{0},\mathbf{P}^{\beta},\mathcal{M})$-generic conditions as follows:
Definition 6.17.
Let $\mathbf{P}$ be a $\varsigma$-iteration of length $\alpha$ and fix an ordinal $\beta<\alpha$. Assume $(\mathbf{P}_{0}^{\beta},\mathbf{P}^{\beta})$ is $\varsigma$-proper. Let $\mathcal{M}\prec H(\theta)$ be countable and such that $\mathbf{P},\beta\in\mathcal{M}$. Let $\{D_{n}:n\in\omega\}$ be a family of open dense subsets of $\mathbf{P}^{\beta}\star[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]\cap\mathcal{M}$. A condition $(p_{0},q_{0})\in\mathbf{P}^{\beta}$ is a $(\mathbf{P}_{0}^{\beta},\mathbf{P}^{\beta},\mathcal{M})$-generic condition for $\{D_{n}:n\in\omega\}$ and $[\mathbf{P}^{(*\beta)}:\mathbf{P}_{0}^{\beta}]$ if there is a family of $\mathbf{P}_{0}^{\beta}$-names $\{\dot{E}_{n}:n\in\omega\}$ such that for each $n\in\omega$, $\circledast_{0}$-$\circledast_{2}$ from Lemma 6.16 hold.
Thus, Lemma 6.16 says that under the given hypothesis, we can always find extensions of conditions in $\mathbf{P}^{\beta}\cap\mathcal{M}$ which are $(\mathbf{P}^{\beta}_{0},\mathbf{P}^{\beta},\mathcal{M})$-generic for the given family of dense open sets and $[\mathbf{P}^{(*\beta)}:\mathbf{P}_{0}^{\beta}]$.
Proof of Lemma 6.16.
The proof splits into two arguments whose proof is quite similar to those of Lemma 6.9 and Lemma 6.10.
Claim. There is a condition $(p^{\prime},q^{\prime})\leq(p,q)$ forcing $D_{n}^{\divideontimes}$ to be an open dense subset of $[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$.
Proof of Claim. First note that $\mathbf{P}^{\beta}$ is strongly proper, so mimicking the proof of 6.9, all we have to do is to define suitable open dense subsets of $\mathbf{P}^{\beta}\cap\mathcal{M}$. This is done as follows:
@@ 3. $\circledast_{2}$
    For any $\mathbf{P}^{\beta}$-name $\dot{r}\in\mathcal{M}$, define $D(\dot{r})$ as,
    $D(\dot{r})=\{(a,b)\in\mathbf{P}^{\beta}\cap\mathcal{M}:(a,b)\Vdash\dot{r}\in[\mathbf{P}^{\beta}:\mathbf{P}^{\beta}_{0}]\vee(a,b)\Vdash\dot{r}\notin[\mathbf{P}^{\beta}:\mathbf{P}^{\beta}_{0}]\}$

### v2

@@ II $\tau_{0}(f\upharpoonright 1)$ $\tau_{0}(f\upharpoonright 2)$ $\tau_{0}(f\upharpoonright 3)$ldots
  5. 4)
    $\Gamma_{<\beta}^{1}=\{\langle q(\beta)[\Gamma_{\beta+1}^{0}]:\beta\in dom(q)\rangle:(\exists p)(p,q)\in\Gamma_{<\beta}\}$.
@@
Let $\mathbf{P}=\langle\mathbb{P}_{\beta},(\dot{\mathbb{Q}}_{\beta}^{0},\dot{\mathbb{Q}}_{\beta}^{1}):\beta<\alpha\rangle$ be a $\varsigma$-iteration and $\beta_{0}\leq\alpha$ some fixed ordinal. Define $\mathbf{P}^{(*\beta_{0})}$ as:
@@
here, the operation $q*\beta_{0}$ keeps the meaning of previous sections and means
$q*\beta_{0}=\{(\gamma,\dot{q}_{\gamma})\in q:\gamma\geq\beta_{0}\}$
The order is given by $(p_{1},q_{1})\leq(p_{0},q_{0})$ if and only if $p_{1}\leq_{\mathbf{P}_{0}}p_{0}$ and for each $\gamma\in dom(q_{0})$, $p_{1}\upharpoonright\gamma\Vdash(p_{1}(\gamma),q_{1}(\gamma))\leq(p_{0}(\gamma),q_{0}(\gamma))$.
Note that for $\beta_{0}=\alpha$, for any $(p,q)\in\mathbf{P}^{(*\beta_{0})}$ we have $q=\emptyset$. If $\Gamma_{<\beta}^{0}$ is a $\mathbf{P}^{\beta}_{0}$-generic filter, define in $V[\Gamma_{<\beta}^{0}]$ the quotient forcing $[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$ and give it the order induced by $\mathbf{P}^{(*\beta)}$.
It can be easily seen that $\mathbf{P}$ can be represented as a $\mathbf{P}^{\beta}_{0}$-restricted iteration of $\mathbf{P}^{\beta}$ and $[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$. Indeed, let
$[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]=\{{((p,q)^{\check{}}},p\upharpoonright\beta):(p,q)\in\mathbf{P}^{(*\beta)}\}$
which is the natural name for the poset $[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$. Now let
$\mathbf{P}^{\beta}*_{\mathbf{P}^{\beta}_{0}}[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]=\{((p,q),{(p_{0},q_{0})^{\check{}}}):(p,q)\in\mathbf{P}^{\beta}\land((p_{0},q_{0})\in\mathbf{P}^{(*\beta)})\land(p\leq p_{0}\upharpoonright\beta)\}$
Note that this poset is a $\mathbf{P}^{\beta}_{0}$-restricted iteration equivalent to $\mathbf{P}$: there is a natural dense embedding $\varphi_{\beta}$ from $\mathbf{P}$ into $\mathbf{P}^{\beta}*_{\mathbf{P}^{\beta}_{0}}[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$ above, to say,
$(p,q)\xrightarrow{\varphi_{\beta}}((p\upharpoonright\beta,q\upharpoonright\beta),(p,q*\beta)^{\check{}})$
For any $((p,q),(p_{0},q_{0})^{\check{}})\in\mathbf{P}^{\beta}*_{\mathbf{P}^{\beta}_{0}}[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$, we have $(p\upharpoonright\beta^{\frown}(p_{0}*\beta),q^{\frown}q_{0})\in\mathbf{P}$, and its $\varphi_{\beta}$-image is below $((p,q),(p_{0},q_{0})^{\check{}})$. The factorization $\mathbf{P}^{\beta}*_{\mathbf{P}_{0}^{\beta}}[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$ will be useful later. Note that if $D\subseteq\mathbf{P}$ is an open dense set, then $\varphi[D]$ gives us a dense subset of $\mathbf{P}^{\beta}*_{\mathbf{P}^{\beta}_{0}}[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$, instead of an open dense subset. This remark comes to say that in the iteration theorem we will prove in the next section, we need to make a convention on how to deal with the just stated remark. If $D\subseteq\mathbf{P}$ is an open dense set and $\beta<\alpha$, let us define the set $\overline{D}^{\beta}$ as follows:
$\overline{D}^{\beta}=\{p\in\mathbf{P}^{\beta}*_{\mathbf{P}^{\beta}_{0}}[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]:(\exists p_{0}\in D)(p\leq\varphi_{\beta}(p_{0})\}$
Similarly, if $\mathcal{M}\prec H(\theta)$ is countable and $\beta,\mathbf{P}\in\mathcal{M}$, for any $D\subseteq\mathbf{P}\cap\mathcal{M}$ open dense, define $\overline{D}^{\beta}$ as
$\overline{D}^{\beta}=\{p\in\mathbf{P}^{\beta}*_{\mathbf{P}^{\beta}_{0}}[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]\cap\mathcal{M}:(\exists p_{0}\in D)(p\leq\varphi_{\beta}(p_{0})\}$
Clearly, if $D\subseteq\mathbf{P}\cap\mathcal{M}$ is open dense, then $\overline{D}^{\beta}$ is an open dense subset of $\mathbf{P}^{\beta}*_{\mathbf{P}^{\beta}_{0}}[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]\cap\mathcal{M}$. This operation will be used only for sets of the form $D\subseteq\mathbf{P}\cap\mathcal{M}$, so there will be no danger of confusion about which of the two versions above is being used (the second one, always). We need one lemma that illustrates the reason of introducing these objects, and the proof of Theorem 8.1 will show how we make use of them. When $D\subseteq\mathbf{P}\cap\mathcal{M}$ and $\Gamma$ is a $\mathbf{P}^{\beta}$-generic filter, define
$D^{\divideontimes}_{\Gamma}=\{(p,q)\in D:(p\upharpoonright\beta,q\upharpoonright\beta)\in\Gamma\}=[D:\mathbf{P}^{\beta}]$
@@
Let $\mathbf{P}$ be a $\varsigma$-iteration of length $\alpha$ and $\mathcal{M}\prec H(\theta)$ countable such that $\mathbf{P}\in\mathcal{M}$ and pick $\beta\in\alpha\cap\mathcal{M}$. Let $D\subseteq\mathbf{P}\cap\mathcal{M}$ be an open dense set and assume $(p,q)\in\mathbf{P}^{\beta}$ is $(\mathbf{P}^{\beta}_{0},\mathbf{P}^{\beta},\mathcal{M})$-generic for $[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$ and $\overline{D}^{\beta}$, and let $\dot{E}$ be a $\mathbf{P}^{\beta}_{0}$-name witnessing this. Then,
$(p,q)\Vdash(\overline{D}^{\beta})^{\divideontimes}=\{(r,s*\beta):(r,s)\in D^{\divideontimes}_{\dot{\Gamma}}\}$
and
$(p,q)\Vdash\dot{E}\subseteq\{(r,s*\beta):(r,s)\in D^{\divideontimes}_{\dot{\Gamma}}\}$
Therefore, $(p,q)$ forces that for any $(r,s)\in(\overline{D}^{\beta})^{\divideontimes}_{\Gamma}$, there is $\tau$ such that $(r,\tau^{\frown}s)\in D$. In particular, this holds for all elements of $\dot{E}_{\Gamma}$.
@@
It is enough to prove the first forcing relation, as the second follows from the first one, and the choice of $(p,q)$ and $\dot{E}$. Let $\Gamma$ be a $\mathbf{P}^{\beta}$-generic filter such that $(p,q)\in\Gamma$. Pick $(r,s)\in(\overline{D}^{\beta})^{\divideontimes}_{\Gamma}$. By definition of $\overline{D}^{\beta}$, there is $(r_{0},s_{0})\in D$ such that:
@@
    $(r_{0}\upharpoonright\beta,s_{0}\upharpoonright\beta)\in\Gamma$.
@@
    $(r,s)\leq(r_{0},s_{0}*\beta)$.
[…]
Convention. Finally, before passing to the next section, whenever we mention a factorization of the form $\mathbf{P}^{\beta}*_{\mathbf{P}^{\beta}_{0}}[\mathbf{P}^{(*\beta)}:\mathbf{P}^{\beta}_{0}]$, we assume this factorization of $\mathbf{P}$ has been done accordingly to the previous paragraphs.
8. A preservation theorem.
We need a few things before going into the preservation theorem. Assume $\mathbf{P}$ is a $\varsigma$-iteration, $\mathcal{M}\prec H(\theta)$ is countable such that $\mathbf{P}\in\mathcal{M}$ and pick a $\mathbf{P}_{0}^{\alpha}$-name $\dot{\mathbb{R}}\in\mathcal{M}$. Given a family $\mathcal{D}=\{D_{n}:n\in\omega\}$ of open dense subsets of $(\mathbf{P}*_{\mathbf{P}_{0}^{\alpha}}\dot{\mathbb{R}})\cap\mathcal{M}$, we say that $\mathcal{D}$ is $\mathbf{P}$-sufficient [^5] It would be more correct to write something like $(\mathbf{P},\mathcal{M})$-sufficient, since this notion depends on $\mathcal{M}$ as well, but to keep simplicity we write $\mathbf{P}$-sufficient. Since this notion is only used in the proof of the next theorem and $\mathcal{M}$ is fixed, there is not danger of confusion. if any $(\mathbf{P}^{\alpha}_{0},\mathbf{P},\mathcal{M})$-generic condition for $\{D_{n}:n\in\omega\}$ and $\dot{\mathbb{R}}$, is also a $(\mathbf{P},\mathcal{M})$-generic condition. If $\{D_{n}:n\in\omega\}$ is the family of dense open subsets of $\mathbf{P}$ living in $\mathcal{M}$, and $E\subseteq(\mathbf{P}*_{\mathbf{P}_{0}^{\alpha}}\dot{\mathbb{R}})\cap\mathcal{M}$ is open dense, it is easy to see that the family $\{H_{n}:n\in\omega\}$ defined as
$H_{n}=\{(p^{\frown}\dot{r},q)\in E:(p,q)\in\bigcap_{j\leq n}D_{j}\}$
is a $\mathbf{P}$-sufficient family of open dense sets. Moreover, for any $\beta\in\alpha\cap\mathcal{M}$, it turns out that $\{\overline{H}_{n}^{\beta}:n\in\omega\}$ is $\mathbf{P}^{\beta}$-sufficient [^6] Let $D\in\mathcal{M}$ be an open dense subset of $\mathbf{P}^{\beta}$ and define $D^{\prime}=\{(p,q)\in D_{0}:(p\upharpoonright\beta,q\upharpoonright\beta)\in D\}$, this is an open dense subset of $\mathbf{P}$ and lives in $\mathcal{M}$, so there is $k\in\omega$ such that $D_{k}=D^{\prime}$. Then we have that for any $((p\upharpoonright\beta,q\upharpoonright\beta),(p^{\frown}\dot{r},q*\beta)^{\check{}})\in\varphi_{\beta}[H_{k}]$, $(p\upharpoonright\beta,q\upharpoonright\beta)\in D$, which implies the same about $\overline{H}_{k}^{\beta}$.. We will denote this family by $\mathcal{H}_{\beta}(E,\mathbf{P},\dot{\mathbb{R}})$, that is $\mathcal{H}_{\beta}(E,\mathbf{P},\dot{\mathbb{R}})=\{\overline{H}_{n}^{\beta}:n\in\omega\}$ and assume this enumeration is fixed.
Theorem 8.1.
Let $\mathbf{P}=\langle\mathbb{P}_{\beta},(\dot{\mathbb{Q}}_{0}^{\beta},\dot{\mathbb{Q}}_{1}^{\beta}):\beta<\alpha\rangle$ be a $\varsigma$-iteration of Cohen proper forcings. Let $\mathcal{M}\prec H(\theta)$ be countable such that $\mathbf{P}\in\mathcal{M}$ and $\dot{\mathbb{R}}\in\mathcal{M}$ a $\mathbf{P}_{0}^{\alpha}$-name for a forcing. Fix $\beta_{0}\in\alpha\cap\mathcal{M}$. Let $D\subseteq(\mathbf{P}*_{\mathbf{P}^{\alpha}_{0}}\dot{\mathbb{R}})\cap\mathcal{M}$ be an open dense set, a condition $(p_{0},q_{0})\in\mathbf{P}^{\beta_{0}}$ and assume that $(p_{0},q_{0})$ is $(\mathbf{P}^{\beta_{0}}_{0},\mathbf{P}^{\beta_{0}},\mathcal{M})$-generic for $[\mathbf{P}^{(*\beta_{0})}*_{\mathbf{P}_{0}^{\alpha}}\dot{\mathbb{R}}:\mathbf{P}^{\beta_{0}}_{0}]$ and $\mathcal{H}_{\beta_{0}}(D,\mathbf{P},\dot{\mathbb{R}})$. Furthermore, assume that $\tilde{\eta}$ is a $\mathbf{P}^{\beta_{0}}_{0}$-name such that $p_{0}\Vdash\tilde{\eta}=(\tilde{\eta}_{0},\tilde{\eta}_{1})\in[\mathbf{P}^{(*\beta_{0})}:\mathbf{P}^{\beta_{0}}_{0}]\cap\mathcal{M}$. Then there is a condition $(\overline{p},\overline{q})\in\mathbf{P}$ and a ${\mathbf{P}}_{0}^{\alpha}$-name $\dot{E}$ such that:
@@
    $p_{0}=\overline{p}\upharpoonright\beta_{0}$ and $q_{0}=\overline{q}\upharpoonright\beta_{0}$.
@@
    $(\overline{p},\overline{q})$ is $(\mathbf{P}_{0}^{\alpha},\mathbf{P},\mathcal{M})$-generic for $[\mathbf{P}^{(*\alpha)}*_{\mathbf{P}_{0}^{\alpha}}\dot{\mathbb{R}}:\mathbf{P}_{0}^{\alpha}](\equiv\dot{\mathbb{R}})$ and $\mathcal{H}_{\alpha}(D,\mathbf{P},\dot{\mathbb{R}})$.
@@
    $\overline{p}\Vdash\dot{E}\subseteq\dot{\mathbb{R}}\cap\mathcal{M}[\dot{\Gamma}_{<\alpha}^{0}]\text{ is open dense}$.
@@
    $(\overline{p},\overline{q})\Vdash\dot{E}\subseteq({D})^{\divideontimes}=\{\dot{r}_{\Gamma}:(\exists(p,q)\in\mathbf{P})((p^{\frown}\dot{r},q)\in D)\}$.